### v2

@@ 7 Conclusion
We have studied the metrological aspects of the Bloch oscillations in single- and many-particle Stark systems which support a phase transition, across the entire spectrum, from an extended to a localized phase. Unlike conventional equilibrium criticality-based quantum sensors, our non-equilibrium probe does not require complex initialization and achieves quantum enhanced sensitivity across the entire extended phase. In addition, our analysis provides an ansatz for the QFI in terms of time, probe size, and number of excitations. In the long-time limit, the QFI scales quadratically with time in both extended and localized phases. In contrast, the precision with respect to the probe size transforms from super-Heisenberg scaling in the extended phase to size-independent behavior in the localized phase. Finally, the number of excitations always enhances the precision of the probe, though, the amount of enhancement depends on the interaction between excitations. We also show that a simple configuration measurement can achieve an estimation accuracy which is not far from the ultimate precision limit. Given its time-dependent feature, our work sets the groundstone for further precision measurement experiments, e.g. in gravitometery or electric field sensing.
@@ Acknowledgements
The authors thank the Center for High Performance Computing at Shanghai Jiao Tong University for providing access to the Siyuan-1 cluster. The work of MA was partially supported by the Yangyang Development Foundation. AB acknowledges support from the National Natural Science Foundation of China (Grants No. 12050410253, No. 92065115, and No. 12274059), and the Ministry of Science and Technology of China (Grant No. QNJ2021167001L). HM acknowledges partial financial support by Iran Optics and Quantum Technologies Development Council.

### v3

@@ 10 Conclusion
We have studied the metrological aspects of the Bloch oscillations in single- and many-particle Stark systems which support a phase transition, across the entire spectrum, from an extended to a localized phase. Unlike conventional equilibrium criticality-based quantum sensors, our non-equilibrium probe does not require complex initialization and achieves quantum-enhanced sensitivity across the entire extended phase. In addition, our analysis provides an ansatz for the QFI in terms of time, probe size, and number of excitations. In the long-time limit, the QFI scales quadratically with time in both extended and localized phases. In contrast, the precision with respect to the probe size transforms from quantum-enhanced scaling in the extended phase to size-independent behavior in the localized phase. This behavior looks very general as, for instance, in the case of multi-particle probes changing the Hamiltonian from non-interacting to interacting case does not change the qualitative behavior of the probe. Interestingly, the number of excitations always enhances the precision of the probe, though, the amount of enhancement depends on the interaction between excitations. We also show that a simple configuration measurement can achieve an estimation accuracy which is not far from the ultimate precision limit. Finally, we have demonstrated a practical scheme in which configuration measurement together with a maximum likelihood estimation approach can closely saturate the Cramér-Rao inequality.
11 Codes and data availability
The computer codes for the simulations of this paper are available at:
https://github.com/hassanmanshouri/The-QFI-of-Stark-model.git
The data used to plot the figures are available upon reasonable requests from the authors.
@@ Acknowledgements
The authors thank the Center for High Performance Computing at Shanghai Jiao Tong University for providing access to the Siyuan-1 cluster. The work of MA was partially supported by the Yangyang Development Foundation. AB acknowledges support from the National Natural Science Foundation of China (Grants No. 12050410253, No. 92065115, and No. 12274059), and the Ministry of Science and Technology of China (Grant No. QNJ2021167001L). HM acknowledges partial financial support by Iran Optics and Quantum Technologies Development Council. HM would like to thank H. Bakhshian for valuable discussions on numerical analysis. This work has been supported by the Center for International Scientific Studies & Collaborations (CISSC), "Ministry of Science, Research and Technology" of Iran.